# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ 2.2 Inexact supervision
Inexact supervision refers to providing coarse-grained labels instead of precise point-wise annotations. The goal is to train a model using rough labeling information to achieve accurate point-level tasks. It typically includes sub-cloud-level annotations [103, 104, 39], scribble annotations [105], bounding box annotations [104, 106, 107, 108, 109, 110, 111, 112], and scene-level annotations [113, 13], among others.
@@
3D semantic segmentation. Liu et al. [104] utilize box-level annotations, as shown in Fig. 4, and combine them with self-training and point activation maps to reduce the cost of point-wise labeling. Unal et al. [105] use scribbles to annotate LiDAR point clouds, as shown in Fig. 5, which also significantly improves the efficiency of point cloud annotation. Some studies [104, 103, 39] use subcloud-level labels to reduce annotation costs, assigning the classes present within the sampled regions as labels. For instance, Wei et al. [103] propose a point cloud scene segmentation network from cloud-level weak labels on indoor 3D data, and they introduce a multipath region mining module. Lin et al. [39] propose a weakly supervised semantic segmentation method for airborne laser scanning (ALS) point clouds. They leverage weak subcloud annotations, as shown in Fig. 6, first generating accurate pseudo-labels using a classification network with overlap region loss and elevation attention, and then optimizing a segmentation network with supervised contrastive loss to efficiently learn from limited subcloud labels. Compared to the above labeling methods, more studies prefer using scene-level labels [113, 53, 114, 40, 13, 115, 116, 98], where only the class information of the entire scene is provided for 3D point clouds, without annotating individual points or object instances, as shown in Fig. 7. For instance, Ren et al. [113] introduce scene-level annotations as supervision while addressing point-level semantic segmentation, 3D proposal generation, and 3D object detection, achieving better results than previous methods. Xia et al. [13] achieve recent performance of fully supervised methods using scene-level annotations propagated to each point and conservative methods for creating pseudo-labels.
@@
In LiDAR Remote Sensing, Wang et al. [40] train by assigning a point to each class label of the sub-cloud, incorporating contextual constraints and pseudo-label generation. They use active learning based on Temporal Output Discrepancy (TOD) to select the most informative sub-point clouds for labeling, achieving near fully-supervised segmentation performance with minimal labeled data on ALS, MLS, and TLS datasets.
3D object detection. Some studies use position-level annotations [118, 89] instead of 3D bounding boxes for 3D object detection to reduce annotation costs. For instance, Meng et al. [118] combine bird’s-eye view (BEV) center-click annotations with a small number of precise 3D instance labels to significantly reduce annotation costs. Scene-level annotation is a simpler approach. WyPR [113] uses simple scene-level labels and addresses the problem of weakly supervised 3D point cloud object detection through multi-task learning and self-supervised constraints. Qin et al. [119] combine scene-level labels with results from a pre-trained network applied to images mapped from 3D point clouds as soft labels, significantly reducing annotation costs while maintaining 3D object detection performance. Wei et al. [120] use 2D bounding box annotations to train a 3D object detector without relying on precise 3D labels.
3D instance segmentation. Some research [106, 108, 107, 109, 110, 111, 112] uses box annotations to reduce labeling costs. For instance, GaPro [107] generates pseudo-labeled masks from 3D bounding boxes to train a 3D point cloud instance segmentation network. Box2Mask [109] replaces dense point-level annotations with 3D bounding box labels, leveraging a voting mechanism and non-maximum clustering to effectively address 3D instance segmentation under weak supervision.
@@ 2.3 Inaccurate supervision
Inaccurate supervision refers to cases where labels contain errors or noise, causing the supervision signal to deviate from the true information, as shown in Fig. 8. Techniques such as sample selection [121, 122] and label correction [123] can mitigate the negative impact of inaccurate labels on model training and improve performance.
Research on robust learning with noisy labels in point clouds is limited. Ye et al. [124] propose the novel Point Noise Adaptive Learning (PNAL) framework, which uses point-by-point confidence of historical predictions for each point to select reliable labels and considers neighborhood correlation to generate the best labels using a voting strategy. Chen et al. [125] consider the paradigm of teacher-student models and design a noise-resistant instance supervision module to eliminate the negative effects of noisy labels from a noise learning perspective. Gao et al. [41] propose a new semantic segmentation framework for handling airborne LiDAR point cloud data with noisy labels, which adaptively corrects noisy labels based on the learning state of different categories to enhance the network’s robustness against noise.
In recent years, 2D foundation models like Contrastive Vision-Language Pre-training (CLIP) [126] and Segment Anything (SAM) [127] have achieved significant success in 2D vision tasks. These models are typically pre-trained on large-scale datasets and then fine-tuned for specific tasks, often outperforming traditional methods. Some studies [128, 129, 130, 131, 132] have applied these powerful 2D foundation models to 3D tasks, enhancing the performance of 3D scene understanding. Pseudo-label. This method uses a foundation model to segment the projected image of the point cloud and then maps the segmentation result back to the point cloud through reprojection or other methods to obtain pseudo-labels [133, 134, 135, 88, 136, 137, 138, 139, 140, 141]. For instance, Segment3D [133] uses high-quality segmentation masks generated by the 2D foundation model SAM to automatically generate labels for point clouds, instead of manual annotations. Dong et al. [134] utilize sparse 2D point annotations as prompts, project 2D semantic segmentation results into 3D space, and combine pseudo-labels from different viewpoints using a voting strategy to generate high-quality 3D semantic pseudo-labels. feature alignment. Chen et al. [142] utilize SAM generates semantic masks from 2D images, which are projected onto 3D point clouds for region-level segmentation and 2D-3D alignment. Dense feature distillation is then applied to transfer 2D features to 3D point clouds.
@@
Continual test-time adaptation - $x^{t}$ No (pre-trained) Yes
@@ 2.4 Domain adaptation
To address this issue, “domain adaptation” aims to transfer knowledge from a labeled source domain to an unlabeled target domain, thereby reducing annotation costs and mitigating the effects of domain shift. It can be regarded as an effective weakly supervised learning approach, as it does not require labeled data from the target domain, offering a practical solution to the challenges of expensive and time-intensive annotation. It includes unsupervised domain adaptation, source free domain adaptation, (continual) test time adaptation, and so on. Their differences are shown in Table 2.
@@ 2.4.1 Unsupervised domain adaptation
Unsupervised Domain Adaptation (UDA) aims to leverage a labeled source domain model to improve performance on an unlabeled target domain. The main challenge is to transfer domain-invariant features from the source to the target domain. Most UDA research has focused on 2D images [143, 144, 145, 146, 144, 147], with only a few studies addressing 3D point clouds. Several approaches have been proposed to tackle the unsupervised domain adaptation problem for 3D point clouds, including adversarial learning, self-training, feature alignment, and so on.
@@
PointDAN [148] is the first work to address point cloud classification in the UDA context, which leverages local and global feature modeling to extract multi-scale information and employs adversarial training to minimize the distribution discrepancy between the source and target domains. Following PointDan, these studies [149, 150, 151, 152, 153] also adopt the adversarial learning strategy to solve the 3D UDA problem. Achituve et al. [154] propose a novel approach for domain adaptation on point clouds by combining the self-supervised task of Deformation Reconstruction (DefRec) with Point Cloud Mixup (PCM). DefRec simulates local deformations of point clouds and reconstructs the original shapes, enabling the model to learn cross-domain geometric and semantic features. PCM enhances generalization by generating new training samples through random sampling and mixing of point clouds. [155, 156] also apply self-supervised learning to UDA tasks.
@@
Self-training is another line for this task, which leverages pseudo-labels to learn target knowledge gradually. Cardace et al. [157] propose the RefRec method, which leverages shape reconstruction tasks to learn global shape descriptors for refining pseudo-labels. Combined with self-training strategies and domain-specific classification heads, the method effectively mitigates domain discrepancies, enhancing the performance of unsupervised domain adaptation for point cloud classification. [158, 159, 160, 161, 162, 163, 164] based self-training mechanisms have achieved SOTA performance in the field of 3D UDA.
Additionally, some methods [165, 166, 167, 168, 169] adopt feature alignment strategies. Wang et al. [169] address the complex domain discrepancies in indoor synthetic-to-real 3D object detection using object-aware augmentation, hierarchical domain alignment (class-level and holistic-level), and pseudo-label optimization strategies.
@@
In recent years, with the rapid development of foundation models, some studies [170, 171, 172] have integrated them to address point cloud UDA tasks. Peng et al. [170] leverage the 2D image embeddings provided by the Segment Anything Model (SAM) as a bridge to align 3D point cloud features from different domains into a unified feature space. Wu et al. [171] utilize the vision-language knowledge of the CLIP model to generate task-specific text embeddings through Visual-driven Prompt Adaptation (VisPA) and achieve multi-modal feature fusion via Text-guided Context Interaction (TexCI). Xu et al. [172] leverage the knowledge priors of Visual Foundation Models (VFMs) to generate high-precision pseudo labels (VFM-PL) and employ a FrustumMixing strategy to semantically mix source and target domain data, reducing distribution discrepancies.
@@
However, research on point cloud UDA in the LiDAR remote sensing domain remains limited. [23, 42, 43, 44, 45, 46] apply the same method: feature alignment. For instance, Luo et al. [23] effectively improve cross-scene MLS point cloud semantic segmentation performance through adaptive data and feature alignment on the target domain point cloud. Peng et al. [42] align local geometric features and global features and reduce the distribution discrepancy between source and target domains through adversarial training. Xie et al. [43] enhance the model generalization through data augmentation techniques and introduce the Deep CORAL method, which aligns second-order statistics of features to achieve cross-city semantic segmentation of ALS point clouds. Shen et al. [44] alleviate the data domain gap by selecting similar scenes based on the scene height distribution, and align the feature distributions using the adversarial learning module (ALM) solving the domain adaptation problem from terrestrial laser scanning (TLS) data to mobile laser scanning (MLS) data.
2.4.2 Source free domain adaptation
@@
Traditional domain adaptation (DA) requires simultaneous access to both source domain data and target domain data for model adaptation. In contrast, source-free domain adaptation (SFDA) [173, 174, 175, 176, 177, 178] relies solely on a pre-trained model from the source domain and performs adaptation to the target domain without access to source domain data. This is achieved by leveraging unlabeled target domain data and utilizing pseudo-labels or feature distributions generated by the model to update its parameters for improved performance in the target domain.
@@
Currently, SFDA is predominantly focused on 2D image processing, while research in 3D point clouds remains relatively limited. [165, 179, 180, 181] apply pseudo-label to solve SFDA. For instance, Saltori et al. [165] propose the first source-free unsupervised domain adaptation framework for LiDAR-based 3D object detection. Wang et al. [180] employ pseudo-label optimization and prototype matching to achieve point cloud primitive segmentation under the source-free domain adaptation paradigm. In contrast to their approach, Tian et al. [182] utilize a Gaussian Mixture Model (GMM) to generate virtual domain samples and enhance the intra-class compactness of the target domain to achieve effective alignment between the target and virtual domains.
2.4.3 Test time adaptation
@@
Test-Time Adaptation (TTA) refers to the process of dynamically optimizing a model during the testing phase using unlabeled target domain data to address domain shift, without accessing source domain data or ground-truth labels. Continual Test-Time Adaptation (CTTA) extends TTA by enabling the model to progressively adapt to continuously changing target domains while retaining its adaptability to previous data distributions, making it suitable for more complex and dynamic scenarios. Most (C)TTA methods are currently applied to 2D images, including techniques such as batch normalization calibration [183, 184], entropy minimization [185, 186], contrastive learning [187, 188], and pseudo-labeling [189, 188].
Research on TTA for 3D point clouds remains relatively limited, which is mainly based on pseudo-labeling technology combined with other techniques [190, 191, 179, 192]. For instance, Cao et al. [191] achieve dynamic adaptation to the target domain through single-modal pseudo-label aggregation, cross-modal pseudo-label fusion, and class-wise momentum queue updates. Saltori et al. [179] enhance pseudo-label coverage through adaptive selection and geometric feature propagation while leveraging temporal consistency regularization to enforce feature smoothness across consecutive frames. Research on TTA is also included in other fields such as point cloud classification [193], registration [194, 195], super-resolution [196], and so on.
For LiDAR remote sensing, Wang et al. [16] propose a Progressive Batch Normalization (PBN) method. PBN incrementally updates BN layers by integrating statistical information from test batches while optimizing learnable BN parameters through information maximization and pseudo-label generation. Experiments are conducted across three adaptation scenarios (e.g., photogrammetric to airborne LiDAR, airborne to mobile LiDAR, and synthetic to mobile LiDAR data) and demonstrate the effectiveness of the method on multiple public datasets.

### v2

@@ 3.2 Inexact supervision
Inexact supervision refers to providing coarse-grained labels instead of precise point-wise annotations. The goal is to train a model using rough labeling information to achieve accurate point-level tasks. It typically includes sub-cloud-level annotations [89], scribble annotations [155], bounding box annotations [156], and scene-level annotations [15], among others.
@@
3D semantic segmentation. Liu et al. [157] utilize box-level annotations, as shown in Fig. 4, and combine them with self-training and point activation maps to reduce the cost of point-wise labeling. Unal et al. [155] use scribbles to annotate LiDAR point clouds, as shown in Fig. 5, which also significantly improves the efficiency of point cloud annotation. Some studies use subcloud-level labels to reduce annotation costs, assigning the classes present within the sampled regions as labels [158, 89]. For instance, Wei et al. [158] propose a point cloud scene segmentation network from cloud-level weak labels on indoor 3D data, and they introduce a multi-path region mining module. Compared to the above labeling methods, more studies prefer using scene-level labels [159, 160, 15], where only the class information of the entire scene is provided for 3D point clouds, without annotating individual points or object instances, as shown in Fig. 6. For instance, Ren et al. [159] extract supervision from scene-level labels within a multi-task framework using multiple-instance learning (MIL) and cross-task consistency losses, demonstrating the feasibility of 3D recognition with only scene labels, but applying scene labels uniformly to all points produces sparse, ambiguous supervision, noisy pseudo-labels, and imprecise boundaries that limit segmentation accuracy. To address this, Xia et al. [15] discover semantic “primitives” in the point-cloud feature space via unsupervised clustering and conservatively assign scene labels to related primitives using bipartite matching, converting global label broadcasting into local associations, thereby densifying supervision and significantly improving pseudo-label quality and segmentation performance.
@@
In LiDAR remote sensing, Lin et al. [89] propose a weakly supervised semantic segmentation method for ALS point clouds. They leverage weak subcloud annotations, as shown in Fig. 7, first generating accurate pseudo-labels using a classification network with overlap region loss and elevation attention, and then optimizing a segmentation network with supervised contrastive loss to efficiently learn from limited subcloud labels. Wang et al. [81] train by assigning a point to each class label of the sub-cloud, incorporating contextual constraints and pseudo-label generation. They use active learning based on Temporal Output Discrepancy (TOD) to select the most informative sub-point clouds for labeling, achieving near fully-supervised segmentation performance with minimal labeled data on ALS, MLS, and TLS datasets. WHCN [90] constructs a hypergraph from high-confidence superpoints derived from scene-level annotations. This hypergraph is then used as input to predict high-precision point-level pseudo-labels, which are subsequently used to train a segmentation network.
3D object detection. Some studies use position-level annotations instead of 3D bounding boxes for 3D object detection to reduce annotation costs [161, 134]. For instance, Meng et al. [161] combine bird’s-eye view (BEV) center-click annotations with a small number of precise 3D instance labels to significantly reduce annotation costs. Using Scene-level annotation is a simpler approach. WyPR [159] uses simple scene-level labels and addresses the problem of weakly supervised 3D point cloud object detection through multi-task learning and self-supervised constraints. Qin et al. [162] combine scene-level labels with results from a pre-trained network applied to images mapped from 3D point clouds as soft labels, significantly reducing annotation costs while maintaining 3D object detection performance.
3D instance segmentation. Some research uses box annotations to reduce labeling costs [163, 164, 156]. For instance, Box2Mask [164] first demonstrates that dense segmentation models can be trained using only bounding boxes: it assigns points to boxes by voting and obtains segmentations via non-maximum suppression, but relies on simple heuristics for overlapping or nested boxes (e.g., assigning points to the smallest-volume box), which induces misassignments and limits the performance ceiling. To address this limitation, GaPro [163] proposes a two-stage paradigm: it first generates high-quality pseudo-labels from bounding boxes and then trains arbitrary segmentation networks on them. GaPro models point membership in overlapping regions probabilistically with Gaussian processes and provides uncertainty estimates, offering greater flexibility and improved performance compared with hard assignments, but it remains constrained by the accuracy of the bounding boxes.
3.2.1 Discussions for inexact supervision
Unlike incomplete supervision, where the primary difficulty lies in propagating sparse labels, inexact supervision faces challenges unique to LiDAR’s modality: coarse labels often span multiple objects or vertical layers, leading to severe label leakage; the geometric boundaries in 3D space are blurred by scene- or subcloud-level supervision; and occlusion effects combined with platform-dependent scanning patterns induce pseudo-label drift. To address these issues, recent methods introduce a dedicated elevation-attention module to enhance critical height information in airborne LiDAR, and employ superpoint generation together with hypergraph convolutional networks to model complex higher-order structural relationships among objects in outdoor scenes. These strategies make weak supervision more aligned with LiDAR’s structural characteristics, enabling more precise segmentation and recognition under coarse labeling conditions.
@@ 3.3 Inaccurate supervision
Inaccurate supervision refers to cases where labels contain errors or noise, causing the supervisory signal to deviate from the true information, as shown in Fig. 8. Techniques like sample selection [165] and label correction [166] help mitigate the negative impact of inaccurate labels on model training and improve performance.
Research on robust learning with noisy labels in point clouds is limited. Ye et al. [167] propose the novel Point Noise Adaptive Learning (PNAL) framework, which uses point-by-point confidence of historical predictions for each point to select reliable labels and considers neighborhood correlation to generate the best labels using a voting strategy. Chen et al. [168] consider the paradigm of teacher-student models and design a noise-resistant instance supervision module to eliminate the negative effects of noisy labels from a noise learning perspective. Gao et al. [91] propose a new semantic segmentation framework for handling ALS point cloud data with noisy labels, which adaptively corrects noisy labels based on the learning state of different categories to enhance network robustness against noise.
In recent years, the significant progress of 2D foundation models, such as CLIP [169] and SAM [170], has provided new opportunities for weakly supervised learning with inaccurate labels. A common approach is to first perform 2D segmentation on the projected images of a point cloud and then maps the segmentation results back to the point cloud through methods like re-projection to generate pseudo-labels, which are used as imprecise annotations [171, 172, 173]. For example, Segment3D [171] utilizes high-quality segmentation masks generated by SAM to automatically label point clouds as a substitute for manual annotations. However, its performance ceiling and generalization capability are entirely constrained by SAM and the domain of the training data.
3.3.1 Discussions for inaccurate supervision
Unlike inexact supervision, where the core difficulty stems from coarse and imprecise annotations, inaccurate supervision in LiDAR exhibits modality-specific failure modes. First, noisy labels tend to propagate structurally through sparse yet extended geometric neighborhoods, amplifying their negative impact. Second, when pseudo-labels are derived from 2D foundation models, cross-modal projection errors and viewpoint inconsistencies exacerbate label noise, especially near object boundaries. Third, LiDAR’s sparse and incomplete geometric structures make it difficult to correct such noise, leading to error accumulation and fragile supervisory signals. Recent works address these challenges through confidence-driven filtering, multi-view voting, and noise-adaptive correction, improving robustness against label noise in large-scale LiDAR segmentation.
@@ 3.3.1 Discussions for inaccurate supervision
Continual test-time adaptation - $x_{1}^{t},x_{2}^{t},......,x_{n}^{t}$ No (pre-trained) Yes
@@ 3.4 Domain adaptation
To address this issue, domain adaptation (DA) aims to transfer knowledge from a labeled source domain to an unlabeled target domain, thereby reducing annotation costs and mitigating the effects of domain shift. It can be regarded as an effective WSL approach, as it does not require labeled data from the target domain, offering a practical solution to the challenges of expensive and time-intensive annotation. It includes unsupervised domain adaptation, source-free domain adaptation, (continual) test-time adaptation, and so on. Their differences are shown in Table 4.
@@ 3.4.1 Unsupervised domain adaptation
Unsupervised Domain Adaptation (UDA) aims to leverage a labeled source domain model to improve performance on an unlabeled target domain. The main challenge is to transfer domain-invariant features from the source to the target domain. Most UDA research has focused on 2D images [174], with only a few studies addressing 3D point clouds. Several approaches have been proposed to tackle the UDA problem for 3D point clouds, including adversarial learning, self-training, feature alignment, and so on.
@@
Pointdan [175] is the first work to address point cloud classification in the UDA context, which leverages local and global feature modeling to extract multi-scale information and employs adversarial training to minimize the distribution discrepancy between the source and target domains. Following Pointdan, these studies also adopt the adversarial learning strategy to solve the 3D UDA problem [176, 177]. Achituve et al. [178] propose a novel approach for DA on point clouds by combining the self-supervised task of Deformation Reconstruction (DefRec) with Point Cloud Mixup (PCM). DefRec simulates local deformations of point clouds and reconstructs the original shapes, enabling the model to learn cross-domain geometric and semantic features. PCM enhances generalization by generating new training samples through random sampling and mixing of point clouds. Others also apply self-supervised learning to UDA tasks [179, 180].
@@
Self-training is another line for this task, which leverages pseudo-labels to learn target knowledge gradually. Cardace et al. [181] propose the RefRec method, which leverages shape reconstruction tasks to learn global shape descriptors for refining pseudo-labels. Combined with self-training strategies and domain-specific classification heads, the method effectively mitigates domain discrepancies, enhancing the performance of UDA for point cloud classification. Some methods based on self-training mechanisms have achieved SOTA performance in 3D UDA [182, 183].
Additionally, some methods adopt feature alignment strategies [184, 185]. Wang et al. [185] address the complex domain discrepancies in indoor synthetic-to-real 3D object detection using object-aware augmentation, hierarchical domain alignment (class-level and holistic-level), and pseudo-label optimization strategies. To address failures of domain adaptation caused by large distribution differences introduced by different LiDAR sensors (e.g., varying beam counts or scanning patterns), SALUDA [186] introduces a self-supervised implicit surface reconstruction auxiliary task that forces the model to reconstruct scene geometry on both source and target data, thereby learning shared geometric features. It strongly depends on geometric information for domain alignment and may be limited when domain gaps are dominated by non-geometric factors (e.g., differing noise patterns).
@@
In recent years, with the rapid development of foundation models, some studies have integrated them to address point cloud UDA tasks[187, 188, 189]. Peng et al. [187] leverage the 2D image embeddings provided by the Segment Anything Model (SAM) as a bridge to align 3D point cloud features from different domains into a unified feature space. Wu et al. [188] utilize the vision-language knowledge of the CLIP model to generate task-specific text embeddings through Visual-driven Prompt Adaptation (VisPA) and achieve multi-modal feature fusion via Text-guided Context Interaction (TexCI). Xu et al. [189] leverage the knowledge priors of Visual Foundation Models (VFMs) to generate high-precision pseudo labels (VFM-PL) and employ a FrustumMixing strategy to semantically mix source and target domain data, reducing distribution discrepancies.
@@
However, research on point cloud UDA in the LiDAR remote sensing domain remains limited. Some methods apply the same strategy [92, 93, 94, 95, 96, 98]: feature alignment. For instance, Luo et al. [92] effectively improve cross-scene MLS point cloud semantic segmentation performance through adaptive data and feature alignment on the target domain point cloud. Peng et al. [93] align local geometric features and global features and reduce the distribution discrepancy between source and target domains through adversarial training. Xie et al. [94] enhance the model generalization through data augmentation techniques and introduce the Deep CORAL method, which aligns second-order statistics of features to achieve cross-city semantic segmentation of ALS point clouds. Shen et al. [95] alleviate the data domain gap by selecting similar scenes based on the scene height distribution, and align the feature distributions using the adversarial learning module (ALM) that solves the DA problem from TLS data to MLS data. Wang et al. [97] integrate a simple geometric transformation data augmentation technique to enable cross-modal learning from LiDAR point clouds to synthetic aperture radar point clouds. PS-UDA [98] employs an L2-paradigm constraint and a Laplace matrix to pre-align cross-scene features. Luo et al. [99]consider reconstructing the continuous potential scene surface of MLS point clouds and transform both the source and target domain MLS point clouds into a canonical domain to achieve point-level distribution alignment.
3.4.2 Source-free domain adaptation
@@
Traditional DA requires simultaneous access to both source domain data and target domain data for model adaptation. In contrast, source-free domain adaptation (SFDA) relies solely on a pre-trained model from the source domain and performs adaptation to the target domain without access to source domain data [190]. This is achieved by leveraging unlabeled target domain data and utilizing pseudo-labels or feature distributions generated by the model to update its parameters for improved performance in the target domain.
@@
Currently, SFDA is predominantly focused on 2D image processing, while research in 3D point clouds remains relatively limited. Some research applies pseudo-labeling to solve SFDA [184, 191, 192]. For instance, Saltori et al. [184] propose the first SFDA framework for LiDAR-based 3D object detection. Wang et al. [192] employ pseudo-label optimization and prototype matching to achieve point cloud primitive segmentation under the SFDA paradigm. Michele et al. [193] introduce a simple reference model and use prediction consistency between it and the model under training as an automatic stopping criterion to mitigate performance degradation caused by training instability during target-domain adaptation. However, its effectiveness partly depends on the quality of the chosen reference model, and its training loss assumes similar class distributions in the source and target domains—an assumption that can break down under large domain shifts.
3.4.3 Test-time adaptation
@@
Test-Time Adaptation (TTA) refers to the process of dynamically optimizing a model during the testing phase using unlabeled target domain data to address domain shift, without accessing source domain data or ground-truth labels. Continual Test-Time Adaptation (CTTA) extends TTA by enabling the model to progressively adapt to continuously changing target domains while retaining its adaptability to previous data distributions, making it suitable for more complex and dynamic scenarios. Most (C)TTA methods are currently applied to 2D images, including techniques such as batch normalization calibration [194], entropy minimization [195], contrastive learning [196], and pseudo-labeling [197].
Research on TTA for 3D point clouds remains relatively limited. The existing methods are mainly based on pseudo-labeling combined with other techniques [198]. For instance, GIPSO [199] pioneers the “source-free online unsupervised domain adaptation (SF-OUDA)” scenario: given only a pretrained model and no source-domain data, it performs online adaptation via geometric feature propagation and self-training, without access to source data or target labels. However, because it relies solely on the LiDAR modality, its performance degrades significantly when LiDAR becomes unreliable (e.g., due to rain, fog, or sensor noise), and its adaptation process lacks explicit mechanisms to prevent catastrophic forgetting, making it hard to balance adapting to new distributions while retaining source-domain knowledge in continuously changing environments. To address these limitations, CoMAC [191] extends the problem to “multi-modal continuous test-time adaptation (MM-CTTA)”. It adaptively shifts reliance to a more reliable modality (e.g., images) when LiDAR quality drops, improving segmentation robustness, and introduces “Class-Wise Momentum Queues” that randomly restore pseudo-source features to “rehearse” previously learned knowledge, thereby mitigating catastrophic forgetting and achieving more stable performance in dynamic environments. Research on TTA is also included in other fields such as point cloud classification [200], registration [201], super-resolution [202], and so on.
For LiDAR remote sensing, Wang et al. [18] are the first to introduce test-time adaptation (TTA) to the geospatial point cloud domain: they progressively update batch-normalization (BN) statistics and fine-tune BN with self-supervised objectives such as information maximization and pseudo-labeling, enabling a pretrained model to adapt to target domains that differ substantially from the source. However, their method follows the conventional TTA paradigm—adapting once to a single static target domain—and does not adequately address catastrophic forgetting in continuously changing data streams. To address this, APCoTTA [100] advances to continuous test-time adaptation (CTTA) and proposes ALS-specific solutions: it uses gradient signals to dynamically select and update “low-confidence” layers while freezing “high-confidence” layers that retain source knowledge; it applies entropy-based confidence filtering to remove unstable samples and reduce erroneous pseudo-labels; and it introduces stochastic parameter interpolation to fuse some updated parameters with the original source parameters. These designs balance adapting to new distributions with preserving prior knowledge and substantially mitigate forgetting and error accumulation. Nonetheless, APCoTTA’s robustness under extreme class imbalance or when encountering novel open-set classes remains to be explored.
3.4.4 Discussions for domain adaption
Unlike inexact or noisy supervision, domain adaptation in LiDAR faces modality-specific challenges arising from spatiotemporal variability and sensor heterogeneity rather than label issues. First, non-stationary spatiotemporal shifts driven by temporal, seasonal, and regional factors lead to dynamic domain drift. Second, cross-platform geometric discrepancies between ALS, TLS, and MLS cause structural distribution breaks, making feature alignment substantially harder. Third, noise pattern and coverage mismatches amplify alignment instability. To address these LiDAR-specific challenges, recent domain-adaptation methods integrate data augmentation, multi-level feature alignment, dynamically updated batch normalization (BN) layers, and other adaptation strategies. Unlike conventional image DA, which mainly deals with static, texture-driven style shifts, LiDAR DA must cope with non-stationary spatiotemporal drift, cross-platform geometric discrepancies, and sparse, irregular sampling, requiring fundamentally different alignment strategies beyond conventional 2D paradigms. However, current research remains relatively limited in scope, with most efforts focusing on cross-dataset adaptation, leaving broader challenges such as temporal domain shift, cross-platform generalization, and multimodal adaptation largely underexplored.
3.5 Domain generalization
Domain generalization (DG) aims to train a model solely on source domain data, enabling it to perform well on unseen target domains—such as new sensors, locations, or weather conditions—without additional training or fine-tuning. Current research is predominantly focused on autonomous driving, which is critical for safety-critical systems. Geometry-driven methods, such as those proposed by Sanchez et al., mitigate sensor discrepancies by constructing computationally intensive “pseudo-dense” point clouds [203]. However, this approach shifts the generalization burden to the localization module at the cost of real-time performance. Kim et al. address this by simulating sparsity during training or by treating a single source domain as a collection of multiple densities to learn density-invariant features [204]. This approach is limited by its excessive focus on density, potentially overlooking differences in object shapes and scene layouts. Li et al. [205] combine density resampling with test-time adaptation to enhance generalization, albeit with a partial loss of inference speed. In a different approach, Lehner et al. [206] employ adversarial learning to generate physically plausible object deformations, thereby improving robustness against anomalous shapes, such as vehicles in accidents. Furthermore, a large-scale empirical study by Eskandar [207] demonstrates that fundamental design choices, such as adjusting anchor box sizes or selecting appropriate training data (e.g., using clear-weather data to generalize to adverse conditions), also significantly impact generalization performance. In contrast, research in the field of LiDAR remote sensing on domain generalization remains relatively scarce.